A solution is 0.10 M Ba(NO3)2 and also 0.10 M AgNO3. Solid KIO3 is added slowly. (a) Calculate the concentration of iodate ion needed to saturate the solution with Ba(IO3)2(s) and with AgIO3(s). (b) Which compound precipitates first? Why?
Answer #1

Given, [Ba(NO3)2] = 0.10 M , [AgNO3] = 0.10

Calculate the molarity of iodate

The Ksp values for Ba(IO3)2 and AgIO3 are :

Ksp for AgIO3 = 3.0 x 10-8 and Ksp for Ba(IO3)2 = 1.57x 10-9

First calculate the iodate concentration needed to saturate the Ba(IO3)2 as shown below:

Ksp = [Ba2+] [IO3-]2

1.57x 10-9 = 0.10 * (2x)2

4x2 = 1.57 x 10-8

so, x = 6.26 x 10-5 M

Thus, [IO3-] =2x = 2 *6.26 x 10-5 M

                       = 1.25 x 10-4

so, 1.25 x 10-4 M iodate concentration needed to saturate the Ba(IO3)2

Now for AgIO3

Ksp = [Ag+] [IO3-]

3.0 x 10-8 = 0.10 M * x

x = 3.0 x 10-7

3.0 x 10-7 M iodate concentration needed to saturate the AgIO3.

b) The Ksp value for Ba(IO3)2 is lower than AgIO3, so Ba(IO3)2 will be first precipitate out.